# Document Valuation in LLM Summaries: A Cluster Shapley Approach

### v2

@@ 7.2 Numerical Comparison Results
In summary, we empirically see that our proposed algorithm, Cluster Shapley, is able to achieve higher accuracy at lower computational costs by leveraging semantic similarity in LLM embeddings of reviews/documents. Unlike other approximation methods, such as Monte Carlo, Truncated Monte Carlo, and Kernel SHAP, which rely on random sampling without utilizing intrinsic semantic properties, Cluster Shapley exploits the semantic similarities encoded in embeddings. By clustering documents based on semantic similarity, our approach achieves more accurate and computationally efficient Shapley value approximations, underscoring the importance of intrinsic semantic information in document valuation. This demonstrates the power of advanced textual representations from LLMs in enhancing document valuation frameworks.
@@ 7.3 Robustness Checks and Extensions
    For completeness, we also compare the approximation error of different algorithms using alternative metrics, including MSE and MAPE, and find that the results are consistent with those shown in Figure 5. See Web Appendix §H.2 for additional details of this robustness check.
@@
    In the main analysis, we use our proposed adaptive DBSCAN (Algorithm 2) to enforce tight distance constraints within clusters. To examine the impact of this design choice, we conduct a robustness check using the standard (non-adaptive) DBSCAN algorithm. We find that while standard DBSCAN performs reasonably well, its performance is consistently inferior to our proposed adaptive version. Details of this comparison are provided in Web Appendix §H.3.
@@
    In the main analysis, we use all 48 test queries and report average performance across them. To assess robustness to query selection and hyperparameter tuning, we conduct a sample-splitting check by randomly dividing the queries into two halves and replicating the analysis for each split. The results, shown in Web Appendix §H.4, confirm consistent comparative performance across splits and stable hyperparameter choices.
@@
    As discussed in §4.4.3, in such large data settings, we can integrate Monte Carlo sampling into our Cluster Shapley to improve efficiency. We present additional experiments in Web Appendix §H.5 demonstrating the performance of our approach for such settings. The first experiment considers a setting where the number of relevant documents $S_{q}=10$. This setup reflects more computationally intensive, yet still feasible, scenarios in which exact Shapley values can be computed. All other settings are consistent with the main analysis. The second experiment involves a setting where the number of relevant documents is extremely large, $S_{q}=30$. This number is unlikely to be commonly used in practice; however, we consider this case as a proof of concept to demonstrate the scalability of our approach. For example, google_citation_count reports that Google AI Overviews include 4–5 citations on average, with a maximum of 33 in rare cases. Since computing exact Shapley values is intractable at this scale, we synthetically construct ground-truth Shapley values and simulate the evaluation function $v(q,A)$. Across both experiments, the Monte Carlo–based Cluster Shapley demonstrates favorable accuracy–cost trade-offs, reinforcing its practicality in realistic and large-scale applications.
@@ 8 Conclusion
Our work makes three core contributions. First, we propose a principled framework for equitable document valuation based on Shapley values, offering a theoretically grounded approach that satisfies key desiderata such as fairness, generality, and scalability. This framework can operate across different summarization methods and evaluation mechanisms, making it broadly applicable across search, review, and Q&A platforms. Second, we introduce Cluster Shapley, a novel approximation algorithm that leverages semantic similarity between documents to reduce the computational burden of traditional Shapley value computation. This method enables a flexible trade-off between efficiency and accuracy via a tunable clustering parameter and is supported by theoretical guarantees. Third, using a real-world case study based on Amazon product reviews, we demonstrate the practical effectiveness of our algorithm: it achieves up to a 40% reduction in computation while keeping the MAPE below 20%, and consistently outperforms existing benchmark methods.
While our proposed framework represents a significant step forward, several important directions remain open for future work. One key avenue is to explore more dynamic or fine-grained valuation mechanisms that account not only for document content but also for metadata such as authorship, credibility, or temporal relevance. Additionally, extensions to multi-turn interactions and conversational query contexts—where user queries evolve over time—could further enrich the valuation framework. Another promising direction is to consider how document value attribution systems can be combined with reputation mechanisms. While a large stream of theoretical and empirical research has established the importance of reputation mechanisms in sustaining two-sided platforms (cripps_etal_2004; resnick_etal_2000; yoganarasimhan_2013), it is unclear how these systems can survive without direct feedback on contributor’s performance. Thus, extending the document valuation framework to augment/replace existing reputation systems can be an important next step. Finally, while Cluster Shapley is an effective first approximation, future work could investigate hybrid methods that combine semantic clustering with adaptive sampling or reinforcement learning to further improve efficiency and attribution fidelity.
@@ Appendix A Illustration Examples of AI-Generated Summaries
The next example, shown in Figure A4, comes from Reddit, a Q&A forum. Reddit is currently piloting a new RAG-enhanced Q&A assistant called Reddit Answers (reddit2025answers). Using the same query, “How to train for a 5K in a month for beginners?”, Reddit Answers generates a summarized response based on user-generated content from relevant Reddit threads. Similar to Google’s AI Overview, the summary is accompanied by links to the original responses, though the content is limited to the Reddit platform.
@@
Figure A5: Amazon AI-Powered Shopping Aisstant - Rufus
Appendix B Prompt Designs
@@ 6. Key Reminders:
Appendix C Temperature Selection in GPT-4o for Summarization and Evaluation Prompts
@@
Detailed results for the summarization and evaluation experiments are presented in Web Appendix C.1 and Web Appendix C.2, respectively. For the summarization, we select a temperature of 0.1 to balance output consistency and expression richness. For the evaluation, we also choose a temperature of 0.1.
C.1 Temperature for Summarization Prompt
Because the outputs of summarization prompts are text information, to gauge the performance, we use two metrics: (1) We assess semantic consistency through embedding-based cosine similarity, which measures the degree to which different summaries preserve the same underlying meaning regardless of specific word choices; (2) We examine lexical diversity using TF-IDF (Term Frequency–Inverse Document Frequency) similarity measures, which quantify the variation in vocabulary and phrasing across multiple generations. While semantic consistency measures reliability in meaning preservation, lexical diversity reflects the model’s creativity in expression. Figure A8 visualizes the effect of temperature settings on both metrics, and Table LABEL:tab:similarity_stats presents the detailed statistical results across all temperature configurations.
@@ C.1 Temperature for Summarization Prompt
C.2 Temperature for Evaluation Prompt
Choosing the temperature for evaluation is straightforward, as the output is a numerical score ranging from 0 to 10. Since our ultimate goal is to obtain stable Shapley values—computed as weighted averages of these scores—we focus directly on how temperature affects the variance of the Shapley values.
@@ C.2 Temperature for Evaluation Prompt
Appendix D Variance Analysis of Shapley Value
@@
We assume that the random summarizations $\{A(S)\}_{S\subseteq D}$ are mutually independent. Given this assumption and independent white noise, and based on the Shapley formula -Equation (4)—which is essentially a weighted average of $v$, the variance in Shapley can be expressed as the weighted average of the variances of $v$. Thus, in the following, we focus on analyzing the variance of $v(A(S))$:
@@
D.1 Empirical Variance in Summarization and Evaluation
To validate this variance decomposition and gauge the magnitudes of these variances, we conduct an experiment to quantify the variance contributions from both the summarization and evaluation stages.
@@ D.1 Empirical Variance in Summarization and Evaluation
  • Total Variance $\text{Var}(v(A(S)))$ — This variance represents the overall variability of the evaluation scores for a given subset, considering all summarization and evaluation rounds. For each subset of reviews, we generate multiple summarizations and perform several evaluation rounds for each summarization. The total variance is calculated as the variance of all the evaluation scores across these rounds, capturing the combined effects of both summarization and evaluation.
  • Summarization Variance $\text{Var}(\mu(A(S)))$ — This variance reflects the variability introduced during the summarization process. After generating multiple summaries for each subset, we compute the mean evaluation score for each summary. The summarization process variance is then calculated as the variance of these mean evaluation scores across different summarizations. This captures how much the content of the summaries themselves contributes to the overall variability in evaluation scores, independent of the evaluation noise.
  • Evaluation Variance $\text{Var}(\varepsilon)$ — This variance isolates the variability introduced during the evaluation process. For each summarization, we evaluate the subset multiple times. The evaluation noise variance is computed as the average variance of the scores within each summarization round, reflecting the inconsistency of GPT-based evaluations across the same summary. In other words, it measures how much the scores fluctuate due to noise in the evaluation model rather than changes in the summaries themselves.
@@
D.2 Reduce the Variance of Shapley Value
@@
To further examine whether evaluations should be replicated three or four times, we conduct an experiment using the test dataset consisting of five queries, each associated with five reviews—the same setup as in Web Appendices C and D.1. We fix the summarization to a single run and vary the number of evaluation repetitions from 1 to 8. For each configuration, we compute Shapley values across 10 independent replications to estimate the variance.
@@
Appendix E Proofs
In this appendix, we present the proofs for all theorems in the main text.
E.1 Proof for Theorem 1
Proof.
$|\phi_{i}-\phi_{j}|\leq L\epsilon.$ (A4)
We now claim that:
$\hat{\phi}_{G_{k}}=\sum_{j\in G_{k}}\phi_{j}.$
$\hat{\phi}_{G_{k}}=\psi(G_{k})=\sum_{j\in G_{k}}\phi_{j}.$
E.2 Proof for Corollary 1
Proof.
E.3 Proof of Theorem 2
Proof.
For any document $i\in G_{k}$, the total error can be decomposed as:
$|\hat{\phi}_{i}-\phi_{i}|\leq L\epsilon.$
Combining both parts completes the error bound:
Finally, regarding computational complexity:
  • The clustering step requires $O(n^{2})$ operations.
Therefore, the total computational complexity is:
$O(n^{2}+C_{\mathcal{A}}(m)).$
E.4 Proof for Corollary 2
Proof.
The total error decomposes as:
where the term $|\hat{\phi}_{i}-\phi_{i}|\leq L\epsilon$.
$\tilde{\phi}_{G_{k}}=\frac{1}{N}\sum_{t=1}^{N}X_{t}.$
By Hoeffding’s inequality, for any $\delta>0$:
To ensure deviation at most $\delta$ with probability $1-\eta$:
$\delta=V_{\max}\sqrt{\frac{\log(2/\eta)}{2N}}.$
Finally, noting $\tilde{\phi}_{i}=\tilde{\phi}_{G_{k}}/|G_{k}|$ gives:
$|\tilde{\phi}_{i}-\hat{\phi}_{i}|\leq\frac{\delta}{|G_{k}|}.$
Combining yields:
Appendix F Empirical Validation of Lipschitz Continuity Assumption
The observed upper bound behavior—where embedding distance provides a ceiling for marginal contribution differences—validates the Lipschitz continuity assumption while capturing the nuanced relationship between semantic similarity and functional contribution in our evaluation framework. This empirical evidence supports the use of embedding-based distances as meaningful constraints in Shapley value computations.
Appendix G Computation time of the Clustering Step

### v3

@@ 7.1 Comparison with Off-the-shelf Shapley Approximation Algorithms
In summary, our empirical results show that the proposed Cluster Shapley algorithm achieves substantially higher accuracy at lower computational costs by leveraging semantic similarity in LLM-based document embeddings. Unlike other approximation methods-such as Monte Carlo, Truncated Monte Carlo, and Kernel SHAP-that rely on random sampling without exploiting intrinsic semantic relationships, Cluster Shapley groups semantically similar documents to approximate cooperative contributions more efficiently. This scalability provides practitioners with a flexible toolkit to balance accuracy and computation cost according to platform constraints. More broadly, these findings highlight the value of incorporating semantic representations from LLMs into document valuation frameworks, enabling practical and principled approaches to fair value attribution at scale.
7.2 Comparison with Simple Attribution Rules
So far, we saw that Cluster Shapley method achieves the best efficiency–accuracy frontier among all Shapley approximation algorithms. Nevertheless, one might argue that simple attribution rules that are not based on Shapley can be both cheap and effective. To that end, we now consider two simple attribution rules – equal attribution and relevance-weighted attribution, and benchmark the performance of our approach against them.
  • $\bullet$
    Equal Attribution: This baseline splits the full-set value uniformly across documents, i.e., all documents in $S_{q}$ get equal attribution $\hat{\phi}^{\mathrm{Equ}}_{i}(q)=v(q,A(q,S_{q}))/|S_{q}|$. This baseline is simple to apply in practice, similar in spirit to Cloudflare’s pay-per-crawl approach (72).
  • $\bullet$
    Relevance-Weighted Attribution: This baseline allocates $v(q,A(q,S_{q}))$ proportionally to the relevance score between the query and each document:
    $s_{i}(q)\;=\;\text{cosine\_similarity}(e_{i},e_{q}),\qquad w_{i}(q)\;=s_{i}(q)/\sum_{j\in S_{q}}s_{j}(q)\$
    $\hat{\phi}^{\mathrm{Rel}}_{i}(q)\;=\;v(q,A(q,S_{q}))\,w_{i}(q),\qquad i\in S_{q}.$
These two simple rules are costless to compute but can introduce substantial bias, as their assigned document values deviate widely from the fair values defined by Shapley. Such comparisons are crucial because platforms must weigh fairness against computational feasibility. In practice, a simple rule such as equal attribution can appear attractive for its negligible cost, as in Cloudflare’s pay-per-crawl model (72), yet understanding how large the fairness loss is essential for responsible design.
Table 3 quantifies this fairness–efficiency trade-off. Panel A reports attribution accuracy relative to the exact Shapley baseline, together with normalized computation cost (where 1 corresponds to the cost of computing all 255 subsets). Panel B displays the corresponding results for Cluster Shapley across different clustering diameters $\epsilon$, listed from coarse to fine (high $\epsilon$ to low), highlighting how accuracy improves as computational investment increases.
The results exhibit a clear pattern: simpler attribution rules incur negligible computational cost but suffer from substantial bias, whereas more accurate methods require greater computation. In Panel A, equal attribution yields an MAPE exceeding 150%, indicating a large deviation from fair allocation, while relevance-weighted attribution achieves a lower but still significant 31.93%. By contrast, even a very coarse Cluster Shapley with $\epsilon=0.70$ achieves a smaller MAPE of 26.49% at an extremely low cost—only 0.02 times that of the exact Shapley (roughly 3.5 out of 255 subset evaluations). Panel B further illustrates the fairness–efficiency trade-off within Cluster Shapley: as $\epsilon$ decreases, computational cost increases while accuracy improves. For instance, with $\epsilon=0.20$, Cluster Shapley achieves an MAPE of 11.85% at 0.60 times the computational cost of the exact Shapley. These findings quantify a tangible “cost of fairness”: achieving fair value attribution demands meaningful, though not prohibitive, computation.
For platform decision-makers, these results provide actionable guidance. When computational budgets are tight, the framework helps quantify how much fairness must be traded off for efficiency. Conversely, when fairness toward content contributors is a design priority, the same analysis clarifies the computational investment required. Although our experiments rely on limited or synthetic query data, the observed bias is likely heterogeneous across query types; once real user-level query distributions become available, the same framework can readily quantify this heterogeneity at scale.
Finally, Cluster Shapley should be viewed as a proof of concept rather than a prescriptive algorithm for deployment. Adapting it to specific LLM-based summarization architectures—or replacing it with improved approximation algorithms—may further enhance the efficiency frontier. The central insight is conceptual: the Shapley-based framework offers a principled way to evaluate and compare the computational cost of fairness in document-value attribution.
Panel A: Accuracy–Efficiency Comparison of Simple Attribution Rules and Exact Shapley
Method MAE MSE MAPE Computation Cost
Exact Shapley 0 0 0 1.00
Equal Attribution 0.3284 0.3955 157.86% 0.00
Relevance-Weighted 0.3374 0.3893 31.93% 0.00
Panel B: Cluster Shapley’s Accuracy–Efficiency Frontier Across $\epsilon$
Cluster Shapley $\epsilon$ MAE MSE MAPE Computation Cost
0.70 0.2305 0.1636 26.49% 0.02
0.60 0.2152 0.1038 24.43% 0.08
0.50 0.1908 0.1074 21.05% 0.18
0.40 0.1972 0.1499 21.35% 0.26
0.30 0.1617 0.1723 17.16% 0.38
0.20 0.0913 0.0441 11.85% 0.60
0.10 0.0507 0.0184 8.62% 0.73
0.01 0.0381 0.0148 7.47% 0.77
Table 3: Approximation error (averaged over all documents) and computation cost of simple attribution rules (Panel A) and Cluster Shapley under varying $\epsilon$ (Panel B). The last column, computation cost, is normalized relative to the exact Shapley, which requires evaluating all 255 unique subsets (assigned value 1). Other methods’ costs are expressed as a fraction of this baseline. Equal and relevance-weighted rules only require $v(q,A(q,S_{q}))$, hence their cost is close to 0 ($1/255\approx 0.00$). Note that when calculating MAPE, we add a small constant (0.1, approximately 10% of the mean Shapley value) to the denominator to prevent near-zero Shapley values from inflating the error.
@@ 7.3 Robustness Checks and Extensions
    For completeness, we also compare the approximation error of different algorithms using alternative metrics, including MSE and MAPE, and find that the results are consistent with those shown in Figure 5. See Web Appendix §D.2 for additional details of this robustness check.
@@
    In the main analysis, we use our proposed adaptive DBSCAN (Algorithm 3) to enforce tight distance constraints within clusters. To examine the impact of this design choice, we conduct a robustness check using the standard (non-adaptive) DBSCAN algorithm. We find that while standard DBSCAN performs reasonably well, its performance is consistently inferior to our proposed adaptive version. Details of this comparison are provided in Web Appendix §D.3.
@@
    In the main analysis, we use all 48 test queries and report average performance across them. To assess robustness to query selection and hyperparameter tuning, we conduct a sample-splitting check by randomly dividing the queries into two halves and replicating the analysis for each split. The results, shown in Web Appendix §D.4, confirm consistent comparative performance across splits and stable hyperparameter choices.
@@
    As discussed in §4.4.3, in such large data settings, we can integrate Monte Carlo sampling into our Cluster Shapley to improve efficiency. We present additional experiments in Web Appendix §D.5 demonstrating the performance of our approach for such settings. The first experiment considers a setting where the number of relevant documents $S_{q}=10$. This setup reflects more computationally intensive, yet still feasible, scenarios in which exact Shapley values can be computed. All other settings are consistent with the main analysis. The second experiment involves a setting where the number of relevant documents is extremely large, $S_{q}=30$. This number is unlikely to be commonly used in practice; however, we consider this case as a proof of concept to demonstrate the scalability of our approach. For example, Gupta (26) reports that Google AI Overviews include 4–5 citations on average, with a maximum of 33 in rare cases. Since computing exact Shapley values is intractable at this scale, we synthetically construct ground-truth Shapley values and simulate the evaluation function $v(q,A)$. Across both experiments, the Monte Carlo–based Cluster Shapley demonstrates favorable accuracy–cost trade-offs, reinforcing its practicality in realistic and large-scale applications.
@@ 8 Conclusion
Our work makes three core contributions. First, we propose a principled framework for fair document valuation based on Shapley values, offering a theoretically grounded approach that satisfies key desiderata such as fairness, generality, and scalability. This framework can operate across different summarization methods and evaluation mechanisms, making it broadly applicable across search, review, and Q&A platforms. Second, we introduce Cluster Shapley, a novel approximation algorithm that leverages semantic similarity between documents to reduce the computational burden of traditional Shapley value computation. This method enables a flexible trade-off between efficiency and accuracy via a tunable clustering parameter and is supported by theoretical guarantees. Third, using a real-world case study based on Amazon product reviews, we demonstrate the practical effectiveness of our algorithm: it achieves up to a 40% reduction in computation while keeping the MAPE below 20%, and consistently outperforms existing benchmark methods.
While our proposed framework represents a significant step forward and provides a proof-of-concept pipeline using a standard RAG system, several important directions remain open for future research. A key avenue is to adapt the framework to real-world LLM summarization systems, which often feature richer retrieval stacks, reranking modules, tool use, and reasoning capabilities. It would also be interesting to explore more dynamic and fine-grained valuation approaches that account not only for document content but also for metadata such as authorship, credibility, and temporal relevance. Moreover, extending the framework to multi-turn interactions and conversational query contexts, where user queries evolve over time, could also be interesting.
Another promising direction is to consider how document value attribution systems can be combined with reputation mechanisms. While a large stream of theoretical and empirical research has established the importance of reputation mechanisms in sustaining two-sided platforms (56; 12; 70), it is unclear how these systems can survive without direct feedback on contributor’s performance. Thus, extending the document valuation framework to augment/replace existing reputation systems can be an important next step.
@@ Appendix A Illustration Examples of AI-Generated Summaries
The next example, shown in Figure A3, comes from Reddit, a Q&A forum. Reddit is currently piloting a new RAG-enhanced Q&A assistant called Reddit Answers (3). Using the same query, “How to train for a 5K in a month for beginners?”, Reddit Answers generates a summarized response based on user-generated content from relevant Reddit threads. Similar to Google’s AI Overview, the summary is accompanied by links to the original responses, though the content is limited to the Reddit platform.
@@
Figure A5: Amazon AI-Powered Shopping Assistant - Rufus
Appendix B Supplementary Materials to Section 4
B.1 Document Clustering with the Distance Constraint
[…]
Algorithm 3 Adaptive Distance-Constrained DBSCAN
  Run the standard DBSCAN with $\texttt{MinPts}=1$:
  • For each document $i$: find $r$-neighborhood $N_{r}(i)=\{j:M_{ij}\leq r\}$.
  • Connect points that are within $r$ distance.
  Check the distance constraint:
end while
B.2 Empirical Validation of Lipschitz Continuity Assumption
The observed upper bound behavior-where embedding distance provides a ceiling for marginal contribution differences-validates the Lipschitz continuity assumption while capturing the nuanced relationship between semantic similarity and functional contribution in our evaluation framework. This empirical evidence supports the use of embedding-based distances as meaningful constraints in Shapley value computations.
B.3 Proof for Theorem 1
Proof.
$|\phi_{i}-\phi_{j}|\leq L\epsilon.$ (A2)
We now claim that:
$\hat{\phi}_{G_{k}}=\sum_{j\in G_{k}}\phi_{j}.$
$\hat{\phi}_{G_{k}}=\psi(G_{k})=\sum_{j\in G_{k}}\phi_{j}.$
B.4 Proof for Corollary 1
Proof.
B.5 Proof of Theorem 2
Proof.
For any document $i\in G_{k}$, the total error can be decomposed as:
$|\hat{\phi}_{i}-\phi_{i}|\leq L\epsilon.$
Combining both parts completes the error bound:
Finally, regarding computational complexity:
  • The clustering step requires $O(n^{2})$ operations.
Therefore, the total computational complexity is:
$O(n^{2}+C_{\mathcal{A}}(m)).$
B.6 Proof for Corollary 2
Proof.
The total error decomposes as:
where the term $|\hat{\phi}_{i}-\phi_{i}|\leq L\epsilon$.
$\tilde{\phi}_{G_{k}}=\frac{1}{N}\sum_{t=1}^{N}X_{t}.$
By Hoeffding’s inequality, for any $\delta>0$:
To ensure deviation at most $\delta$ with probability $1-\eta$:
$\delta=V_{\max}\sqrt{\frac{\log(2/\eta)}{2N}}.$
Finally, noting $\tilde{\phi}_{i}=\tilde{\phi}_{G_{k}}/|G_{k}|$ gives:
$|\tilde{\phi}_{i}-\hat{\phi}_{i}|\leq\frac{\delta}{|G_{k}|}.$
Combining yields:
Appendix C Supplementary Implementation Details
C.1 Introduction to RAG
AI search engines are designed to provide real-time, contextually relevant responses to user queries. A key technique behind many of these systems is RAG, which integrates pre-trained LLMs with information retrieval to enhance response accuracy and relevance (2). RAG addresses the limitations of static, pre-trained LLMs by incorporating new information from up-to-date, domain-relevant documents (which can be potentially proprietary to the firm). By grounding responses in reliable documents, RAG improves the relevance of AI-generated answers, reduces hallucinations, and mitigates the issue of outdated information that plagues static models (1).
  • $\bullet$
  • $\bullet$
  • $\bullet$
C.2 Prompt Designs
@@ 6. Key Reminders:
C.3 Temperature Selection in GPT-4o for Summarization and Evaluation Prompts
@@
Because the outputs of summarization prompts are text information, to gauge the performance, we use two metrics: (1) We assess semantic consistency through embedding-based cosine similarity, which measures the degree to which different summaries preserve the same underlying meaning regardless of specific word choices; (2) We examine lexical diversity using TF-IDF (Term Frequency–Inverse Document Frequency) similarity measures, which quantify the variation in vocabulary and phrasing across multiple generations. While semantic consistency measures reliability in meaning preservation, lexical diversity reflects the model’s creativity in expression. Figure A9 visualizes the effect of temperature settings on both metrics.
@@
Choosing the temperature for evaluation is similar and straightforward, as the output is a numerical score ranging from 0 to 10. Since our ultimate goal is to obtain stable Shapley values, which are computed as weighted averages of these scores, we focus directly on how temperature affects the variance of the Shapley values.
@@
C.4 Variance Analysis of Summarization, Evaluation, and Shapley Value
@@
We assume that the random summarizations $\{A(S)\}_{S\subseteq D}$ are mutually independent. Given this assumption and independent white noise, and based on the Shapley formula -Equation (4)-which is essentially a weighted average of $v$, the variance in Shapley can be expressed as the weighted average of the variances of $v$. Thus, in the following, we focus on analyzing the variance of $v(A(S))$:
@@
C.4.1 Empirical Variance in Summarization and Evaluation
@@
  • Total Variance $\text{Var}(v(A(S)))$ - This variance represents the overall variability of the evaluation scores for a given subset, considering all summarization and evaluation rounds. For each subset of reviews, we generate multiple summarizations and perform several evaluation rounds for each summarization. The total variance is calculated as the variance of all the evaluation scores across these rounds, capturing the combined effects of both summarization and evaluation.
  • Summarization Variance $\text{Var}(\mu(A(S)))$ - This variance reflects the variability introduced during the summarization process. After generating multiple summaries for each subset, we compute the mean evaluation score for each summary. The summarization process variance is then calculated as the variance of these mean evaluation scores across different summarizations. This captures how much the content of the summaries themselves contributes to the overall variability in evaluation scores, independent of the evaluation noise.
  • Evaluation Variance $\text{Var}(\varepsilon)$ - This variance isolates the variability introduced during the evaluation process. For each summarization, we evaluate the subset multiple times. The evaluation noise variance is computed as the average variance of the scores within each summarization round, reflecting the inconsistency of GPT-based evaluations across the same summary. In other words, it measures how much the scores fluctuate due to noise in the evaluation model rather than changes in the summaries themselves.
@@
C.4.2 Reduce the Variance of Shapley Value
@@
To further examine whether evaluations should be replicated three or four times, we conduct an experiment using the test dataset consisting of five queries, each associated with five reviews-the same setup as in Web Appendices C.3 and C.4.1. We fix the summarization to a single run and vary the number of evaluation repetitions from 1 to 8. For each configuration, we compute Shapley values across 10 independent replications to estimate the variance.
@@
C.5 Computation time of the Clustering Step